An airplane flies from Pittsburgh to harrisburg in 1.5 hours and returns in 2 hours. If the wind speed is 16 miles per hour blowing from the west, find the air speed of the plane
Answer by ewatrrr(24785)

 
Hi,
An airplane flies from Pittsburgh to harrisburg in 1.5 hours and returns in 2 hours.
Pittburgh basically due west from harrisburg(with the wind going)
Let (s +16) and (s-16) represent the rate to and from respectively
Question states*** Note: d = r*t
(s + 16mph)*1.5 = (s-16mph)*2
1.5s + 1.5*16 = 2s - 2*16
3.5*16 = .5s
56 = .5s
s = 56/.5 = 112mph
CHECKING our Answer***
1.5hr * 128mph = 2hr*96mph = 192mi (distance between Pittsburgh and Harrisburg)